Given f(x) = 3 x+8 and g(x) = sqrt(x+3), find f ° g(4) and g ° f(4), giving exact answers (using sqrt if necessary).

Answer by stanbon(75887)

Given f(x) = 3 x+8 and g(x) = sqrt(x+3), find f ° g(4) and g ° f(4), giving exact answers (using sqrt if necessary).
-------------------
fog(4) = f[g(4)] = f[sqrt(4+3)]
=f[sqrt(7)]
= 3sqrt(7)+8
==============
gof(4) = g[f(4)]
=g[3*4+8]
=g(20)
=sqrt(20+3)
=sqrt(23)
=================
Cheers,
Stan H.